# Leveraging the Power of Large Language Models in Entity Linking via Adaptive Routing and Targeted Reasoning

### v1

@@ 1 Introduction
Traditional EL methods depend on large labeled datasets and extensive fine-tuning, which limits their adaptability to new domains and their ability to scale Ayoola et al. (2022); Wu et al. (2020b). They also struggle with “hard cases” including: 1) low-context mentions, which have insufficient context to disambiguate (e.g., “Ireland” in a short snippet), 2) lexical ambiguity, when the same surface form maps to multiple entity types or senses (e.g., a country versus its national team), and 3) knowledge-intensive cases, which demand external or implicit world knowledge for correct linking (e.g., interpreting ‘the Big Apple’ as New York City involves background knowledge rather than textual clues).
Large Language Models (LLMs) have transformed entity linking by using prompt-based methods (Xin et al., 2024) in place of large annotated datasets thus improving performance on difficult disambiguation cases through their advanced reasoning. However, current LLM-centric approaches are inefficient, as they require repeated model calls at each step, leading to increased cost and slower processing.
@@
ARTER is a structured pipeline as illustrated in Figure  1. The pipeline starts with generating candidate entities and computing candidate scores; then using these scoring signals, a lightweight router classifies mentions as easy or hard cases. Easy mentions are handled by fast traditional EL models, while hard cases are routed through a reasoning prompt for refined disambiguation. This design offers two key benefits. First, it avoids the computational cost of full-model fine-tuning required by systems like ReFinED, which typically need to be adapted to new datasets or domains to achieve strong performance. In our approach, ReFinED remains frozen, serving for fast high-recall candidate generation and handling “easy” mentions. Second, by applying targeted reasoning prompts to difficult mentions, our method avoids relying solely on prompt-based inference as in OneNet Liu et al. (2024) and is more efficient as only hard cases incur the LLM cost rather than every mention. We evaluate ARTER on standard benchmarks—AIDA Hoffart et al. (2011), MSNBC Cucerzan (2007), ACE2004 Ratinov et al. (2011), AQUAINT Milne and Witten (2008), CWEB (Gabrilovich et al., 2013), and WIKI Guo and Barbosa (2018) and demonstrate improved performance through efficient, selective use of LLM reasoning. Based on our benchmarking results, we try to answer the following research questions:
@@ 2 Related Work
Entity Linking (EL) has seen significant progress with systems like BLINK and ReFinED. Wu et al. (2020a) proposed BLINK, a zero-shot framework that uses a bi-encoder for candidate retrieval and a cross-encoder for ranking. While effective on large datasets such as TACKBP-2010, BLINK struggles with long-tail entities due to limited contextual understanding. ReFinED (Ayoola et al., 2022) addresses efficiency by processing all mentions in a single pass and improves accuracy with fine-grained entity typing and entity priors. Our work builds on these systems by using its candidate generation module to obtain high-quality entity candidates, while avoiding the burden of fine-tuning by relying on prompt-based strategies instead.
Despite our focus on EL, recent progress in Relation Extraction (RE) is closely related because both tasks require understanding and reasoning over entity–context relationships. Many prompting techniques developed for RE, such as few-shot and chain-of-thought reasoning, have proven transferable to EL and help LLMs perform contextual disambiguation. Recent work explores LLM-based approaches to both EL and RE. Wadhwa et al. (2021) show that few-shot prompting can match RE baselines, while Chain-of-Thought (CoT) improves accuracy. Wang et al. (2024) apply retrieval-augmentation for entity disambiguation, but find performance still lags behind SOTA on GERBIL (Hoffart et al., 2015). LLMAEL (Xin et al., 2024) augments mention context with LLM-generated descriptions but requires fine-tuning. OneNet (Liu et al., 2024) uses a fully prompt-based EL pipeline without retraining, yet lacks adaptive control over reasoning depth.
@@ 3.4.1 Prompting Strategies
To fully utilize the reasoning capabilities of LLMs, we experimented with various prompting strategies, including zero-shot (Kojima et al., 2023), few-shot (Brown et al., 2020), Chain-of-Thought (CoT) (Wei et al., 2023), contrastive prompting (Chia et al., 2023), and ReAct-style prompting (Yao et al., 2023). Contrastive prompting enhances robustness by providing positive and negative examples, while ReAct-style prompting interleaves reasoning with actions, enabling step-by-step thinking and tool or information access. A detailed comparison of all strategies on the ACE2004 hard cases dataset is shown in Table 8. The compasison shows that the most effective approach turns out to be the one combining few-shot examples with CoT reasoning. This method encourages the model to analyze mentions and candidates step by step, leveraging entity descriptions and category consistency. The prompt format explicitly aligns each mention and its surrounding context with a list of candidate entities and requires the model to identify the most semantically compatible option or return 1-1 if none fits. Few-shot examples illustrate the step-by-step reasoning process that links contextual cues to the appropriate entity description, grounding predictions in the candidate information provided.
@@ 4.1 Datasets
A comprehensive and diverse evaluation was conducted on publicly available datasets. These datasets cover a wide range of domains by including both standard benchmarks and challenging cases, like long-tail entity linking. The datasets used to benchmark performance are the following: AIDA Hoffart et al. (2011), MSNBC Cucerzan (2007), ACE2004 Ratinov et al. (2011), AQUAINT Milne and Witten (2008), CWEB (Gabrilovich et al., 2013) and WIKI Guo and Barbosa (2018). Table 2 mentions details about the datasets. Standard pre-processing steps—such as text normalization, entity extraction, and format standardization were applied for consistency across datasets. Wikipedia dumps were used as the knowledge base for candidate generation, providing an up-to-date and extensive entity source.
@@ 4.3 Entity Linking Baselines
As our primary external baseline, we compared with the state-of-the-art ReFinED model Ayoola et al. (2022). After routing, mentions classified as easy are resolved by ReFinED, which leverages fine-grained entity types and description-based ranking to disambiguate entities without additional reasoning. For secondary analysis, we included the full prompting baseline, which applies LLM-based reasoning to every mention without routing.
@@ 6 Deployment Efficiency
Modern entity linking systems, such as OneNet Liu et al. (2024), use LLMs with reasoning strategies like Chain-of-Thought for every mention. However, not all mentions require this level of reasoning. Our method introduces a router that identifies "easy" mentions and handles them with a lightweight model like ReFinED, skipping LLM reasoning when unnecessary.

### v2

@@ 1 Introduction
Traditional EL methods depend on large labeled datasets and extensive fine-tuning, which limits their adaptability to new domains and their ability to scale ayoola-etal-2022-refined; wu2019zero. They also struggle with “hard cases” including: 1) low-context mentions, which have insufficient context to disambiguate (e.g., “Ireland” in a short snippet), 2) lexical ambiguity, when the same surface form maps to multiple entity types or senses (e.g., a country versus its national team), and 3) knowledge-intensive cases, which demand external or implicit world knowledge for correct linking (e.g., interpreting ‘the Big Apple’ as New York City involves background knowledge rather than textual clues).
Large Language Models (LLMs) have transformed entity linking by using prompt-based methods (xin2024llmaellargelanguagemodels) in place of large annotated datasets thus improving performance on difficult disambiguation cases through their advanced reasoning. However, current LLM-centric approaches are inefficient, as they require repeated model calls at each step, leading to increased cost and slower processing.
@@
ARTER is a structured pipeline as illustrated in Figure  1. The pipeline starts with generating candidate entities and computing candidate scores; then using these scoring signals, a lightweight router classifies mentions as easy or hard cases. Easy mentions are handled by fast traditional EL models, while hard cases are routed through a reasoning prompt for refined disambiguation. This design offers two key benefits. First, it avoids the computational cost of full-model fine-tuning required by systems like ReFinED, which typically need to be adapted to new datasets or domains to achieve strong performance. In our approach, ReFinED remains frozen, serving for fast high-recall candidate generation and handling “easy” mentions. Second, by applying targeted reasoning prompts to difficult mentions, our method avoids relying solely on prompt-based inference as in OneNet liu-etal-2024-onenet and is more efficient as only hard cases incur the LLM cost rather than every mention. We evaluate ARTER on standard benchmarks—AIDA hoffart-etal-2011-robust, MSNBC cucerzan-2007-large, ACE2004 ratinov-etal-2011-local, AQUAINT milne2008learning, CWEB (gabrilovich2013facc1), and WIKI guo2018robust_preprint and demonstrate improved performance through efficient, selective use of LLM reasoning. Based on our benchmarking results, we try to answer the following research questions:
@@ 2 Related Work
Entity Linking (EL) has seen significant progress with systems like BLINK and ReFinED. wu-etal-2020-scalable proposed BLINK, a zero-shot framework that uses a bi-encoder for candidate retrieval and a cross-encoder for ranking. While effective on large datasets such as TACKBP-2010, BLINK struggles with long-tail entities due to limited contextual understanding. ReFinED (ayoola-etal-2022-refined) addresses efficiency by processing all mentions in a single pass and improves accuracy with fine-grained entity typing and entity priors. Our work builds on these systems by using its candidate generation module to obtain high-quality entity candidates, while avoiding the burden of fine-tuning by relying on prompt-based strategies instead.
Despite our focus on EL, recent progress in Relation Extraction (RE) is closely related because both tasks require understanding and reasoning over entity–context relationships. Many prompting techniques developed for RE, such as few-shot and chain-of-thought reasoning, have proven transferable to EL and help LLMs perform contextual disambiguation. Recent work explores LLM-based approaches to both EL and RE. wadhwa2021revisiting show that few-shot prompting can match RE baselines, while Chain-of-Thought (CoT) improves accuracy. entity-disambiguation-fusion-decoding apply retrieval-augmentation for entity disambiguation, but find performance still lags behind SOTA on GERBIL (hoffart2015gerbil). LLMAEL (xin2024llmaellargelanguagemodels) augments mention context with LLM-generated descriptions but requires fine-tuning. OneNet (liu-etal-2024-onenet) uses a fully prompt-based EL pipeline without retraining, yet lacks adaptive control over reasoning depth.
@@ 3.4.1 Prompting Strategies
To fully utilize the reasoning capabilities of LLMs, we experimented with various prompting strategies, including zero-shot (kojima2023zeroshot), few-shot (brown2020fewshot), Chain-of-Thought (CoT) (wei2023cot), contrastive prompting (chia2023contrastive), and ReAct-style prompting (yao2023reactsynergizingreasoningacting). Contrastive prompting enhances robustness by providing positive and negative examples, while ReAct-style prompting interleaves reasoning with actions, enabling step-by-step thinking and tool or information access. A detailed comparison of all strategies on the ACE2004 hard cases dataset is shown in Table 8. The compasison shows that the most effective approach turns out to be the one combining few-shot examples with CoT reasoning. This method encourages the model to analyze mentions and candidates step by step, leveraging entity descriptions and category consistency. The prompt format explicitly aligns each mention and its surrounding context with a list of candidate entities and requires the model to identify the most semantically compatible option or return 1-1 if none fits. Few-shot examples illustrate the step-by-step reasoning process that links contextual cues to the appropriate entity description, grounding predictions in the candidate information provided.
@@ 4.1 Datasets
A comprehensive and diverse evaluation was conducted on publicly available datasets. These datasets cover a wide range of domains by including both standard benchmarks and challenging cases, like long-tail entity linking. The datasets used to benchmark performance are the following: AIDA hoffart-etal-2011-robust, MSNBC cucerzan-2007-large, ACE2004 ratinov-etal-2011-local, AQUAINT milne2008learning, CWEB (gabrilovich2013facc1) and WIKI guo2018robust_preprint. Table 2 mentions details about the datasets. Standard pre-processing steps—such as text normalization, entity extraction, and format standardization were applied for consistency across datasets. Wikipedia dumps were used as the knowledge base for candidate generation, providing an up-to-date and extensive entity source.
@@ 4.3 Entity Linking Baselines
As our primary external baseline, we compared with the state-of-the-art ReFinED model ayoola-etal-2022-refined. After routing, mentions classified as easy are resolved by ReFinED, which leverages fine-grained entity types and description-based ranking to disambiguate entities without additional reasoning. For secondary analysis, we included the full prompting baseline, which applies LLM-based reasoning to every mention without routing.
@@ 6 Deployment Efficiency
Modern entity linking systems, such as OneNet liu-etal-2024-onenet, use LLMs with reasoning strategies like Chain-of-Thought for every mention. However, not all mentions require this level of reasoning. Our method introduces a router that identifies "easy" mentions and handles them with a lightweight model like ReFinED, skipping LLM reasoning when unnecessary.